A solution containing 8.32 g of para-aminobenzoic acid (PABA) and 150 g of chloroform boils at 62.62℃. At the same pressure, pure chloroform boils at 61.15℃. What is the molecular mass of PABA? The Kfor chloroform is 3.63℃/m.


1
Expert's answer
2020-10-07T07:33:06-0400

The empirical formula for PABA"= C_7H_2NO_2" , with a molar mass of "137.14g\/mol"

Chloroform"CHCl_3" has a molar mass of "119.38g\/mol"

"\therefore" "W_{C_7H_{2}NO_2}" "=8.2g"


"W_{CHCl_3}=150g"


"\Delta _{{T}_{b}}" "=61.15\u00b0C"

"m=" "W_{C_7H_2NO_2}\over mC_7H_2NO_2" "\times" "1000\over w"

"m=" "8.2\over 137.14" "\times" "1000\over 1000" "=0.0598"

"0.62\over 3.63\times 0.0598" "\times" "1000" "=134.60"

Molecular mass of PABA"=134.60g\/mol"